# LiveScene: Language Embedding Interactive Radiance Fields for Physical Scene Rendering and Control

### v2

@@ 1 Introduction
Interactive objects are prevalent in our daily lives, and modeling interactable scenes from the real physical world plays an essential role in various research fields, including content generation [33, 35, 52], animation [58, 37, 30], virtual reality [50, 9, 32], robotics [1, 61, 44], and world understanding [16, 14, 53]. This paper tackles the challenging and rarely explored task of reconstructing and controlling multiple interactive objects in complex scenes from a single, casually captured monocular video without previous independent modeling of geometry and kinematics. Prior research on interactable scene modeling, such as CoNeRF [20] and K-Planes [10], typically adopts a joint modeling approach, combining spatial coordinates and all interaction variables as input and representing interactive scene by either implicit MLPs or feature planes. Meanwhile, CoGS [63] learns parameter offsets for different scene parts using multiple independent MLPs after establishing a 3D deformable Gaussian scene. However, these methods primarily focus on capturing interactions for a single object within a clear background, such as a single drawer, toy car, or face [66, 20, 63, 68]. As modeling extends from single objects to multiple objects in complex scenes, as shown in Figure. 1, the interaction spaces become increasingly high-dimensional, complicating these methods for accurate modeling and significantly increasing computational time and memory cost, e.g., 4$\times$ A100 GPU for 2 weeks to converge training in CoNeRF [20] and 500M Gaussian storage for a regular indoor living room in CoGS [63]. Moreover, natural language is an intuitive and necessary interface for interacting with 3D scenes, but language embedding of interactive scenes faces an even more daunting challenge: interaction variation inconsistency. For instance, methods like LERF [23], and OpenNeRF [67], which distill CLIP features into static 3D fields, suffer from significant failures when confronted with scene topology structure changes induced by interactions, such as the distinct structures variation of a cabinet before and after opening.

### v3

@@ Abstract
Abstract
This supplementary material accompanies the main paper by providing more details for reproducibility as well as additional evaluations and qualitative results to to verify the effectiveness and robustness of LiveScene:
$\triangleright$ Section. 7: Configurations of OmniSim and InterReal dataset, including scene assets, interaction variables generation, mask and prompts annotation, and dataset visualization.
$\triangleright$ Section. 8: Video demonstration and anonymous link: https://livescenes. github.io.
$\triangleright$ Section. 9: Additional implementation details.
$\triangleright$ Section. 10: Additional experimental results, including more ablation studies, detailed view synthesis quality comparison, interactive scenes geometry comparison and language grounding comparison.
@@ 1 Introduction
Interactive objects are prevalent in our daily lives, and modeling interactable scenes from the real physical world plays an essential role in various research fields, including content generation [34, 37, 55], animation [61, 39, 31], virtual reality [53, 8, 33], robotics [47, 73, 74, 35, 46], and world understanding [20, 14, 56, 19, 11]. This paper tackles the challenging and rarely explored task of reconstructing and controlling multiple interactive objects in complex scenes from a single, casually captured monocular video without previous independent modeling of geometry and kinematics. Prior research on interactable scene modeling, such as CoNeRF [21] and K-Planes [9], typically adopts a joint modeling approach, combining spatial coordinates and all interaction variables as input and representing interactive scene by either implicit MLPs or feature planes. Meanwhile, CoGS [65] learns parameter offsets for different scene parts using multiple independent MLPs after establishing a 3D deformable Gaussian scene. However, these methods primarily focus on capturing interactions for a single object within a clear background, such as a single drawer, toy car, or face [68, 21, 65, 70]. As modeling extends from single objects to multiple objects in complex scenes, as shown in Figure. 1, the interaction spaces become increasingly high-dimensional, complicating these methods for accurate modeling and significantly increasing computational time and memory cost, e.g., 4$\times$ A100 GPU for 2 weeks to converge training in CoNeRF [21] and 500M Gaussian storage for a regular indoor living room in CoGS [65]. Moreover, natural language is an intuitive and necessary interface for interacting with 3D scenes, but language embedding of interactive scenes faces an even more daunting challenge: interaction variation inconsistency. For instance, methods like LERF [24], and OpenNeRF [69], which distill CLIP features into static 3D fields, suffer from significant failures when confronted with scene topology structure changes induced by interactions, such as the distinct structures variation of a cabinet before and after opening.